OpenStudy (anonymous):

Please help - step by step so I can comprend; 5/8=11/4x+2 - Thanks

hero (hero):

\[\frac{5}{8} = \frac{11}{4x+2}\] Cross multiply to get: 5(4x+2) = 8*11 Distribute to get 5*4x + 2*5 = 8*11 Multiply that to get: 20x + 10 = 88 Subtract 10 from both sides to get 20x = 78 Divide both sides by 20 to get x = 78/20

OpenStudy (anonymous):

Thanks-

hero (hero):

You still have to reduce it